OpenStudy (anonymous):

Hello! How do I solve the differential equation: y''+y=1/(cos(x)) Thanks in advance!

OpenStudy (anonymous):

In more detail: How can I solve the particular integral 1/cos x ?

OpenStudy (zarkon):

\[\frac{1}{\cos(x)}=\sec{x}\] \[\int\sec(x)dx=\int\sec(x)\frac{\sec(x)+\tan(x)}{\sec(x)+\tan(x)}dx\] \[\int\frac{\sec^2(x)+\sec(x)\tan(x)}{\sec(x)+\tan(x)}dx\] let \[u=\sec(x)+\tan(x)\] answer should be easy to get from there

OpenStudy (anonymous):

Thank you for the integral, however I do not need solely integrate 1/cos(x), I have to find a trial function y for which y''+y=1/cos (x) The homogeneous solution is easy, given roots of the auxiliary equation i and -i, it is: y=Acos(x)+Bsin(x) The general solution is then y=Acos(x)+Bsin(x)+PI and PI is then the function y, which satisfies y''+y=1/cos(x)

OpenStudy (zarkon):

Use variation of parameters technique

OpenStudy (anonymous):

Thanks, found a good tutorial for that technique and I think I might find the solution with that